14 pairwise different positive integers are written on the board. Their mean value is equal to 19. Let M be the largest of these numbers. Find the smallest possible value of M.


1
Expert's answer
2021-10-11T03:01:59-0400

given

mean value of these integers=19

total number of integers=14

mean value=sum of numbers"\div" total number of integers

19=sum of numbers"\div" 14

sum of numbers="19\times 24"

As we sum in A.P

"s_n=\frac{n}{2}[2a+(n-1)d]\\n=14\\s_n=19\times 14\\d=1\space (unique \space difference)"

"19\times14=\frac{14}{2}[2a+(14-1)1]\\19=\frac{1}{2}[2a+13]\\38=2a+13\\2a=25\\a=12.5
 \ is smallest possible value of M"